Gdeinward:

I have a factoring question: x^4 − 17x^2 + 16. I need some assistance.

DCharger96:

step 1: Factor x4−17x2+16 x4−17x2+16 =(x+1)(x−1)(x+4)(x−4)

Gdeinward:

Any old chum can find the answer. I need the steps.

DCharger96:

step 2: yo answer (x+1)(x−1)(x+4)(x−4)

DCharger96:

Since both terms are perfect squares, factor using the difference of squares formula, a 2 − b 2 = ( a + b ) ( a − b ) where a = x and b = 4 . ( x + 4 ) ( x − 4 ) ( x + 1 ) ( x − 1 )

DCharger96:

does that help you at all bro

Hero:

Let \(u = x^2\)

Hero:

Then the expression becomes \(u^2 - 17u + 16\)

Hero:

@Gdeinward

Gdeinward:

What method is that? I am unfamiliar with it

Hero:

u-substitution

Hero:

Technically, you don't have to use it, however, it enables you to factor the expression easier. You can put the \(x^2\) back after factoring.

Gdeinward:

I always enjoy easier methods, but I dont know what the teacher will think, seeing a method not introduced in the lesson. I did find how to solve it in that way, but I was worried about it. What are your thoughts? Should I play it safe, or proceed with the u-substitution method?

Hero:

Proceed with the u-substitution. Just say you learned it trying to figure out how to factor it. No biggie. He or she can't dock points for you trying to learn. It's praiseworthy to try to figure out how to do something. Not something that is frowned upon.

Gdeinward:

That is true, I had those same thoughts and assumptions, but you can understand my concerns.

Gdeinward:

On that note, I can take it the rest of the way. Thanks for your assistance.

Hero:

Trust me, you needn't worry no matter what the teacher says or thinks. Its a method that will help you later on in your math studies. If the teacher can't appreciate that, then he or she does not have your best interests in mind.

Gdeinward:

Thank you Hero and DC for your assistance.

DCharger96:

your welcome bro